# HERA: High-efficiency Matrix Compression via Element Replacement

### v1

@@ 2.1 Basic algorithm
8:end for
9:for each pair $x_{pair_{i,j}}$ in $X_{\text{pairs}}$ do
10:     if $x_{pair_{i,j}}[1]<x_{pair_{i,j}}[2]$ then
11:         $X[iter]_{small_{i,j}}\leftarrow x_{pair_{i,j}}[1]$
12:         $X[iter]_{big_{i,j}}\leftarrow x_{pair_{i,j}}[2]$
13:         $fm[iter]_{i,j}\leftarrow 1$
14:     else
15:         $X[iter]_{small_{i,j}}\leftarrow x_{pair_{i,j}}[2]$
16:         $X[iter]_{big_{i,j}}\leftarrow x_{pair_{i,j}}[1]$
17:         $fm[iter]_{i,j}\leftarrow 0$
18:     end if
19:end for
20:for $X_{small}$ and $X_{large}$ do
21:     code[$iter$], codebook [$iter$] $\leftarrow$ PQ($X[iter]_{small}$)
22:     code[$iter$], codebook[$iter$] $\leftarrow$ PQ($X[iter]_{large}$)
23:end for
24:$iter$leftarrow $iter$ 1
25:if $iter$neqthen
26:     Goto line 1
27:end if
However, these transformations require specific parameters for each column and involve complex calculations, which can be computationally intensive and demand significant storage resources. To address this, we employ a reordering method that permutations the elements of the matrix based on their relative sizes. This approach achieves a more uniform distribution, enabling more effective application of compression operations.
Algorithm 3 Dequantization of Product Quantization with HERA heuristics
1:Quantized feature matrices: $code_{small}$, $code_{big}$, codebooks: $codebook_{small}$, $codebook_{big}$, feature maps: $fm$, number of iterations in HERA: $iter$
2:Reconstructed matrix: $\hat{X}\in\mathbb{R}^{N\times D}$
3:Initialize $\hat{X}_{pairs}\in\mathbb{R}^{N\times D/2}$
4:for each iteration from $1$ to $iter$ do
5:     for each pair $x_{pair_{i,j}}$ in $\hat{X}_{pairs}$ do
6:         $\hat{X}_{small_{i,j}}\leftarrow\text{decode}(code_{small}[iter],codebook_{small}[iter])$
7:         $\hat{X}_{big_{i,j}}\leftarrow\text{decode}(code_{big}[iter],codebook_{big}[iter])$
8:     end for
9:     for each pair $x_{pair_{i,j}}$ in $\hat{X}_{pairs}$ do
10:         if $fm[iter]_{i,j}=1$ then
11:              $x_{pair_{i,j}}[1]\leftarrow\hat{X}_{small_{i,j}}$
12:              $x_{pair_{i,j}}[2]\leftarrow\hat{X}_{big_{i,j}}$
13:         else
14:              $x_{pair_{i,j}}[2]\leftarrow\hat{X}_{small_{i,j}}$
15:              $x_{pair_{i,j}}[1]\leftarrow\hat{X}_{big_{i,j}}$
16:         end if
17:     end for
18:end for
19:for each dimension $j$ do
20:     for each element $i$ in $\hat{X}_{pairs}$ do
21:         $\hat{x}_{2i,j}\leftarrow x_{pair_{i,j}}[1]$
22:         $\hat{x}_{2i+1,j}\leftarrow x_{pair_{i,j}}[2]$
23:     end for
24:end for
The purpose of permutation is to explore all possible reordering methods within the search space. For an $N\times D$ matrix, there are $(N!)^{D}$ possible permutations, creating a vast search space.
To manage this, we employ heuristic algorithms to reduce the search space and optimize the algorithm’s complexity. Heuristic algorithms provide approximate solutions by focusing on the most promising areas of the search space, rather than exhaustively evaluating all possible permutations. These methods significantly decrease computational requirements while still delivering high-quality solutions.
By adopting this heuristic reordering strategy, we aim to achieve a more balanced distribution of data across the matrix. This not only simplifies the preprocessing steps but also enhances the efficiency of subsequent clustering and compression processes. Ultimately, this method helps overcome the challenges posed by uneven data distributions, leading to more accurate and resource-efficient clustering outcomes.
2.2 HERA heuristics
To improve quantization and avoid uneven distribution, HERA first perceives the distribution of data sizes, then rearranges the elements according to the distribution, and finally quantizes the rearranged matrix.
2.2.1 HERA Design
We observe that compressing matrix elements of similar sizes as a single column (group) yields better results.HERA uses a low-cost method to locally adjust element arrangement based on data sizes. We first introude the HERA design. The design of HERA heuristics is shown in Figure  4.
Figure 4: HERA heuristics.
HERA groups every two elements in the matrix together, rearranges the matrix elements based on the size of adjacent elements, and then quantizes the rearranged results.
Quantization: The quantization process is illustrated in the algorithm 2, the quantization process consists of multiple steps, which involve rearranging the input data, comparing elements, and quantizing both smaller and larger value sets.
The algorithm begins by rearranging the input matrix $X\in\mathbb{R}^{N\times D}$ into pairs (lines 1 in Algorithm 2). Each element $x_{i}\in X$ is grouped into $X_{\text{pairs}}\in\mathbb{R}^{N\times D/2}$ by combining every two consecutive elements along each dimension. For each dimension $j$, the pairs are formed as $(x_{2i,j},x_{2i+1,j})$ (lines 2–4 in Algorithm 2).
Next, each pair $x_{pair_{i,j}}$ is compared to separate the smaller and larger values. If the first element of the pair is smaller, it is assigned to $X_{small}$ and the second to $X_{big}$, with a feature map $fm$ recording the comparison result as 1 (lines 5–8 in Algorithm 2). Otherwise, the assignment is reversed, and the feature map records a 0 (lines 9–12 in Algorithm 2).
The quantization process proceeds by applying the PQ algorithm to both $X_{small}$ and $X_{big}$. This process is iteratively repeated for a predefined number of iterations ($iter$), refining the quantization results with each iteration (lines 13–16 in Algorithm 2).
Dequantization:The Decompression process consists of multiple steps, namely decoding the quantized values, reconstructing the pairs, and restoring the original matrix dimensions.
The algorithm begins by initializing the reconstructed matrix $\hat{X}_{pairs}\in\mathbb{R}^{N\times D/2}$ (line 1 in Algorithm 3). For each iteration, the algorithm decodes the quantized values from the codebooks for both small and large pairs (lines 2–4 in Algorithm 3). The decoded values are stored in $\hat{X}_{small}$ and $\hat{X}_{big}$ respectively.
Next, for each pair in $\hat{X}_{pairs}$, the algorithm reconstructs the pairs based on the feature map $fm[iter]$. If the feature map value is 1, the first element of the pair is assigned the decoded small value and the second element the decoded big value (lines 5–8 in Algorithm 3). If the feature map value is 0, the assignment is reversed (lines 9–12 in Algorithm 3).
Finally, the pairs are expanded back into the original dimensions to form the reconstructed matrix $\hat{X}$ (lines 13–17 in Algorithm 3).
@@ 3.1 Experiment setup
Dataset: The datasets used in our experiments were generated with matrices of size $N\times D$, where $N$ is the number of samples and $D$ is the dimensionality of each sample. We employed one types of normal distribution to create these datasets.
Normal Distribution Dataset: For the normal distribution dataset, each element of the matrix was drawn from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16. The distribution was truncated to the open interval $(0,1)$ to ensure all values remain within this range.
$\mathbf{X}_{ij}\sim\mathcal{TN}(0.5,0.16,0,1),\quad\forall i\in\{1,\ldots,N\},\quad\forall j\in\{1,\ldots,D\}$
Here, $\mathcal{TN}(\mu,\sigma,a,b)$ denotes the truncated normal distribution with mean $\mu$, standard deviation $\sigma$, truncated to the interval $(a,b)$.
@@ 3.2 Experiment result
We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  7-  7, Figures  10-  10, Figures  13-  13, Figures  16-  16 respectively. For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
@@
Parameter sensitivity analysis: We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance. The results for 1 to 4 iterations are shown in Figures  1919, Figures  $2$, and Figures  2525, respectively. Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).

### v2

@@ 3.2.1 HETA Dequantization
8:Step 2: Reverse Subspace Grouping
9:Combine subspaces $\{G_{i}^{\prime}\}$, where $i=1,2,\dots,m$, $G_{i}^{\prime}=\{\mathbf{g_{i}^{\prime}}\}$ for $i=1$ to $n$, to form $X^{\prime*}$
10:Step 3: Reverse Recursive Partitioning
11:$k\leftarrow l$
12:while $k\neq 0$ do
13:     for $i=1$ to $2^{k-1}$ do
14:         Merge $S^{i}_{k}$ and $L^{i}_{k}$ based on $I_{k}$ and $X^{\prime*}$, then update the results into $X^{\prime*}$
15:     end for
16:     $k\leftarrow k-1$
17:end while
18:Return Reconstructed matrix $X^{\prime}$
3.3 Optimizations
In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO). RO takes advantage of the fact that, after applying HETA, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE. On the other hand, BPO utilizes batch processing techniques to enhance processing speed. Since the codebook used in the HETA algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
Table 1: Range Ratios, and Codebook(CB) Proportions in different compression ratios.
Compression Ratio 4.0 8.0 12.0 16.0
Range Ratio (%) 1.6 4.0 9.2 29.7
CB Proportion (%) 87.5 78.1 71.8 62.5
3.3.1 Residual Optimization
Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the HETA algorithm. As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range. Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Optimization (RO). The RO algorithm, as depicted in Figure 1, begins with the quantization of the original matrix to produce a quantized matrix. This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix. The residual matrix typically exhibits a significantly reduced data range, making it easier to compress. This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or HETA.
During the dequantization process, the quantized residual is dequantized, and this dequantized residual is added back to the initially dequantized matrix to produce a revised dequantized matrix.
Figure 1: Residual Optimization Algorithm.
3.3.2 Batch Processing Optimization
The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the HETA algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
In the original process, elements of the matrix are compared and sorted individually, leading to a sequential, element-by-element approach that can become computationally intensive as the matrix size increases. This method is not optimal for large-scale matrices.
BPO introduces a more advanced technique by processing matrix elements in parallel batches rather than sequentially. This parallelization allows the algorithm to handle multiple comparisons and sorting operations simultaneously, rather than one at a time. By doing so, it reduces the overall computational complexity and accelerates the entire quantization process.
3.3.3 Codebook Quantization Optimization (CQO)
The codebook occupies a significant portion of the space in the HETA algorithm. As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space. Therefore, compressing the codebook can significantly reduce the space used by the algorithm. Our optimization mainly employs the RTN method to compress the codebook.
4 Theory
@@ 5.1 Experiment setup
Dataset: To evaluate the effectiveness of our proposed algorithm, we conducted experiments using one synthetic dataset and four real-world datasets: a synthetic normal distribution dataset, LLM weight dataset, KV cache dataset, Gist dataset, and Wiki dataset. Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text. The results for the other datasets are provided in the appendix (Section A.1). Below, we provide a detailed description of each dataset.
(1) Synthetic normal distribution dataset: This dataset was generated by drawing each element of the matrix from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16. Additionally, one out of every ten thousand elements was replaced with an outlier value, randomly chosen between -100 and 100.
(2) LLM weight dataset: This dataset comprises weight matrices extracted from the large language model (LLM) LLaMA2 (Touvron et al. (2023)).
(3) KV cache dataset: The KV cache dataset is derived from the key-value pairs stored in the cache during inference in KV Quant (Hooper et al. (2024)). The key and value are stored in matrices, respectively.
(4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks. This dataset is particularly useful for evaluating the algorithm’s effectiveness in handling high-dimensional and dense data representations. (Jegou et al. (2010)) (Oliva & Torralba (2001))
(5) Wiki dataset: The Wiki dataset contains document embeddings derived from Wikipedia articles. This dataset tests the algorithm’s ability to compress and process large-scale text data efficiently.
@@
We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  4-  4, Figures  7-  7, Figures  10-  10, Figures  13-  13 respectively. For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
@@
We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance. The results for 1 to 4 iterations are shown in Figures  1616, Figures  1919, and Figures  $2$, respectively. Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).
@@ 6 Conclusion
Appendix A Appendix
A.1 Experimental Results on Additional Datasets